# SADA: Semantic adversarial unsupervised domain adaptation for Temporal Action Localization

### v2

@@ Abstract
Temporal Action Localization (TAL) is a complex task that poses relevant challenges, particularly when attempting to generalize on new – unseen – domains in real-world applications. These scenarios, despite realistic, are often neglected in the literature, exposing these solutions to important performance degradation. In this work, we tackle this issue by introducing, for the first time, an approach for Unsupervised Domain Adaptation (UDA) in sparse TAL, which we refer to as Semantic Adversarial unsupervised Domain Adaptation (SADA). Our contributions are threefold: (1) we pioneer the development of a domain adaptation model that operates on realistic sparse action detection benchmarks; (2) we tackle the limitations of global-distribution alignment techniques by introducing a novel adversarial loss that is sensitive to local class distributions, ensuring finer-grained adaptation; and (3) we present a novel set of benchmarks based on EpicKitchens100 and CharadesEgo, that evaluate multiple domain shifts in a comprehensive manner. Our experiments indicate that SADA improves the adaptation across domains when compared to fully supervised state-of-the-art and alternative UDA methods, attaining a performance boost of up to mAP.
Index Terms:
Artificial Intelligence, Computer Vision, Video Understanding, Domain Adaptation1 Introduction
Recent advances in the field of video understanding have played a critical role in the surge of novel video-based applications –e.g., video indexing, summarization, recommendation, or surveillance. A critical task of this field is Temporal Action Localization (TAL), which involves identifying actions in a video consisting of both their time intervals and action categories. This is particularly difficult given the inherent variabilities of videos. These can be presented, among others, in the form of appearance variability –e.g., different kitchens and/or lighting conditions–, acquisition variability –e.g., different recording devices– or viewpoint variability –e.g., first- or third-person. All in all, these prompt a certain degree of confusion between similar actions to be discriminated.
@@
Unsupervised Domain Adaptation (UDA) has recently become a hot topic given its potential to leverage unlabelled data to mitigate this domain-induced degradation [3, 4, 5]. Unfortunately, despite its considerable success in image-based tasks [6], the application of UDA to video understanding remains underexplored. In fact, to the best of our knowledge, no prior work addresses UDA for TAL setups. The closest proposal is SSTDA [7], which approaches the problem of action segmentation. This focuses on making per-frame action predictions (i.e., dense TAL), tackling datasets where no concurrent actions take place [8, 9]. This approach inevitably requires additional smoothing techniques to preserve temporal coherence. This motivates the focus of this paper on sparse TAL –i.e., segment-level predictions– avoiding the need for additional losses, while intrinsically adapting to multi-label scenarios.
@@
Consequently, in this paper, we propose the first Unsupervised Domain Adaptation method for sparse multi-label detection on TAL, which we name Semantic Adversarial unsupervised Domain Adaptation or SADA for short. Our goal is to minimize the discrepancy between feature representations of a labeled source domain and an unlabelled target domain. These features are extracted with a multi-resolution architecture [10] that we couple with a novel adversarial loss that improves the limitations of existing UDA works. Concretely, existing works normally align domain distributions globally [3]. We propose instead to use pseudo-labeling [11] to assign an action or background (no action) class to each feature representation. With this, our loss factorizes the global adaptation strategy into independent per-class and background alignments – i.e., aligning each action’s distribution across both domains. This results in a more sensitive alignment strategy compared to a global distribution approach, less prone to feature misalignment and, as we will show, better performing.
@@
Assessing the effectiveness of UDA methods for video understanding is a challenging, still unresolved task. Existing proposals on action segmentation [7] follow a subject-based strategy where they aim to adapt a model to new unseen subjects. Here we refer to subject as a person appearing in a video. Nevertheless, little data is normally available from a single subject, which inevitably requires grouping several of them for training. This allows the model to generalize over the subject variability under study, making it unsuitable for domain adaptation. To address these limitations, we draw inspiration from the work of [12] on action recognition and investigate the impact of viewpoint domain shifts on sparse TAL in CharadesEgo [13]. However, we contend that a more comprehensive evaluation necessitates setups with more controllable shifts. For this, we also propose a suite of 6 new setups based on EpicKitchens100 [14] which study the effect appearance and acquisition domain shifts. These comprehensive benchmarks demonstrate that SADA mitigates the performance degradation, improving by a large margin the existing fully supervised (namely source-only) and UDA-based proposals. In short, our main contributions are:
@@
We propose for the first time an unsupervised domain adaptation method suitable for sparse detection scenarios on TAL.
@@ 2 Related work
Temporal Action Localization. At the time of this writing, most of the literature on the task of Temporal Action Localization follows a traditional source-only approach. In other words, they restrict the models’ visibility solely to a training domain, while other domains seen during testing are not available. These works can be categorized as follows: (1) Anchor-based methods [15, 16, 17, 18, 19, 20, 21, 10] propose a two-stage pipeline, consisting of a proposal generation and classification. The first applies heuristic methods – e.g., uniform sampling [20, 15, 21, 10] or action boundaries’ grouping [22, 23] – to generate a dense set of proposals – i.e., temporal segments. In the second stage, they leverage a learnable classifier to predict the corresponding action class and localization offsets of every anchor. Our work falls into this category motivated by the recent success of these methods achieving state-of-the-art results in many TAL benchmarks [24, 25, 14] (2) Anchor-free methods [26, 27, 17, 28] avoid this two-stage approach making per-frame predictions of their corresponding action labels. These methods, however, often suffer from a tendency towards over-segmentation given the potential discrepancy between neighboring frames. Consequently, they require often complex smoothing techniques to improve the boundary predictions [7] (3) Query-based methods [29, 30, 31] recently emerged as an alternative paradigm that follows the principles presented by [32]. This approach exploits the use of a Transformer encoder-decoder architecture [33] to learn a fixed small set of queries given refined video features, each identifying one potential action segment. Intuitively, this results in a non-heuristic-based proposal generation. This comes with the limitation of an increased rigidity, as the number of proposals needs to be fixed beforehand.
@@
Unsupervised Domain Adaptation. Domain Adaptation techniques emerge as an effective solution to bridge the gap between data collected from a source and a target distribution, respectively. A large suite of approaches has been proposed to perform this alignment between labeled and unlabeled domains – e.g., discrepancy minimization [34, 35], entropy minimization [36, 37] or contrastive learning [38]. The most popular approach is arguably the use of adversarial training methods [3, 4, 39, 40, 41, 42, 5], which incorporate a domain classifier trained to discern if samples come from the source or the target domain, in a min-max fashion. This results in the computation of domain-invariant embeddings [43]. Despite convenient, the simplicity of these methods might degrade the quality of the alignment [11], as they potentially align embeddings of source and target domain that represent different semantic information – e.g., different class labels. Few works have been proposed to do this alignment in a more sensitive way [11]. We highlight the importance of [2], proposing to minimize instead the distance of per-class centroids, computed as the mean of the set of predicted feature embeddings of a given class. Its effectiveness, however, relies on the assumption that the data is distributed somewhat homogeneously around the center, as otherwise, the centroids are not necessarily meaningful. In our work, we couple the advantages of both adversarial domain adaptation and semantic alignment and propose for the first time a pure adversarial semantic loss that yields domain invariant representations in a semantically meaningful way, without making explicit assumptions of the distributions (see Fig. 1).
@@
Unsupervised Domain Adaptation for TAL: Despite the considerable success of UDA methods, their applicability has been mostly restricted to image-based scenarios [6] such as image classification [44, 3, 45, 4, 46], object detection [47, 48] or semantic segmentation [49, 50]. Much less attention has been dedicated to video-based applications such as action recognition [51, 52, 53] or spatio-temporal action segmentation [54, 55]. To the best of our knowledge, at the time of this writing, there is no direct comparison with our work focusing on UDA for sparse Temporal Action Localization. The closest work is SSTDA [7] that applies unsupervised domain adaptation for Action Segmentation. SSTDA proposes the use of two global-distribution-based auxiliary tasks to jointly align cross-domain feature spaces. Unlike our proposal, their work falls into the category of anchor-free, making per-frame action predictions. This restricts its applicability to action segmentation scenarios, where current datasets [8, 9] are designed to deal with frame-based single-action classification. In our work, we overcome this limitation by leveraging an anchor-based architecture that enables a natural adaptation to more realistic multi-label scenarios.
@@ 3.1 Problem definition and notation
In this paper, we address the problem of unsupervised domain adaptation for TAL. For this, we define a source domain and a target domain . Domain consists of labeled input videos , where each video is a sequence of frames with . Here contains the begin, end, and class actions of the ground-truth (GT) segments of the video, respectively. The target domain is similar to but lacks the GT information. Concretely, it consists of unlabeled input videos . Our goal is to train a model that can identify the action segments, including both segment coordinates and action labels, in videos from domain , while minimizing the performance degradation on the unlabeled domain .
@@ 3.2 Framework overview
In this work, we propose a model based on a feature pyramid and a classification and localization head (see Fig. 2). This architecture is coupled with a novel semantic adversarial loss that aligns the embeddings of the source and target domain in a semantically meaningful way. More in detail, the model takes as input two videos from domain and , respectively. The model first processes the raw input videos using a frozen pre-trained video backbone. The resulting embeddings of both domains are then passed through a shared SGP pyramid [10] that outputs a set of multi-resolution anchor embeddings for each of the predefined resolution levels. The main goal of our model is to make these anchor embeddings domain invariant. For this, we introduce a level-wise semantic adversarial loss that learns in an adversarial manner to align the embeddings of both domains belonging to a given action class at a given resolution level . Recall that GT information is only available for domain , therefore we rely on the use of pseudo labeling techniques [11] to infer the probable class labels of the data from domain . Finally, we use the domain invariant anchors of domain to train a classification and localization head that learns the underlying tasks in a standard supervised fashion. In short, this permits to learn a classification and localization head that minimizes the decline of performance when applied to the unseen domain .
@@ 3.3 Backbone and SGP pyramid
Our model first takes two input videos and of both domains and . For simplicity, both videos and have length , which we enforce using padding. The method then processes the two videos applying a frozen pre-trained backbone – e.g., I3D [56], Slowfast [57]. This permits to extract, in an effective way, temporal cues of the video into a set of refined video features. These embeddings are then fed to an SGP feature pyramid [10] which combines the use of SGP blocks and the progressive downsampling of the temporal length by a ratio of . This outputs a set of multi-resolution anchor embeddings and , for the two domains, respectively. Here denotes the set of predefined resolution levels and the anchor embeddings of level of a given domain. Concretely, this permits to obtain embeddings for a set of uniformly sampled anchors at each of the resolution levels. The use of a multi-resolution model is favorable to naturally adapt to different action lengths.
@@ 3.4 Classification and localization head
To learn the underlying TAL task, we train in a fully supervised manner a classification and localization module with the labeled source domain . Due to the anchor-based nature of our model, we first require a matching strategy between the set of candidate anchors to the actual GT segments. For this, we follow a center sampling strategy [58, 59, 21, 10]. In other words, for a given level , we define an anchor as positive if the time instant that it represents is near the center of an action. All the rest are marked as negative. We define and as the begins, ends, and action classes of their matching GT segments. We identify negative anchors by assigning them to action class 0.
With this, we design a classification head that maps each of the anchor embeddings to their class distribution. More specifically, we model this as a sequence of 1D convolutions, and train it using a sigmoid focal loss (SFL) [60] as follows:
Similarly, we model a localization head identically as , which predicts the begin-end offsets. We thus define the localization loss as a standard mean squared error (MSE) loss over the positive samples only:
where refers to the positive samples from the -th level only and to the concatenation operation. This yields the final task loss defined as:
@@ 3.5 Our proposal: Semantic adversarial multi-resolution alignment
One of the main contributions of this paper is the design of a novel adversarial-based loss that we name SADA loss, which attempts to overcome the limitations of the extensively used global adversarial loss [3]. Traditional adversarial domain adaptation relies on the idea of designing a domain classifier that learns to identify the domain that each of the embeddings belongs to. The rest of the model learns concurrently the opposite objective, which, as shown by [3], results in the learning of the domain invariant representations. While this approach has been shown to be effective in other fields – e.g., image classification, and object detection – we find that its performance in a more challenging setup like TAL still presents important challenges. One of the main issues, as argued by [11] is that this loss often suffers from feature misalignment which greatly declines the model’s performance. This refers to the cases where these methods align embeddings of non-matching class labels – e.g., aligning embeddings of domain of an action with embeddings of domain of a class .
@@
Local adversarial alignment: To fix this feature misalignment in realistic scenarios like TAL, we propose an alternative adversarial loss formulation that provides a finer-grained alignment. This loss first attempts to perform a local class-aware alignment. This is, for every given resolution level , we group the anchor embeddings of and corresponding to an action label . This is straightforward for the source domain as we have the GT information. For the target domain, we use a hard-pseudo labeling strategy [11] that classifies a given embedding as class if this is the highest-confidence score of the predicted class distribution, and this is above a threshold . Formally we define the hard pseudo-label of a given anchor as:
@@
where are the predicted class probabilities of the anchors. We assign a class to the background anchors – those not assigned to any class. From this, we obtain the new grouped embeddings of source and target domain of class on level as:
@@
for , . Also, is the GT action label of anchor and is its computed pseudo-label from Eq. 4. We then adversarially train a domain classifier to identify the domain of each of these embeddings using a binary cross entropy (BCE) loss:
@@
where and are the labels of each of the domains. Following [3], we introduce a Reverse Gradient Layer (GRL) before the discriminator to invert the sign of the gradient which creates the min-max game where the feature extractor learns to confuse the discriminator. We condition the discriminator to class using a learnable class embedding that we replicate for every selected anchor into an embedding .
@@
Local and global alignment (SADA): Eq. 7 aligns solely the anchor embeddings that are classified as one of the classes, but what happens with the background embeddings that fall below the threshold ? In this case, the loss ignores their influence, yielding a suboptimal partial alignment (see Table IV).
@@
To overcome this issue, we propose our final SADA loss which attempts to combine the best of both global alignment loss [3] and Eq. 7. For this, we introduce a new loss term for the background embeddings as follows:
@@
where again and are the selected background anchors, and is the learnable background embedding. Coupling Eq. 7 and Eq. 8 yields the final formulation of our proposed loss, combining class-wise alignment with the background anchors alignment. Formally:
@@
where is a hyper-parameter that regulates the importance of level on the final alignment loss.

### v3

@@ Abstract
Temporal Action Localization (TAL) is a complex task that poses relevant challenges, particularly when attempting to generalize on new – unseen – domains in real-world applications. These scenarios, despite realistic, are often neglected in the literature, exposing these solutions to important performance degradation. In this work, we tackle this issue by introducing, for the first time, an approach for Unsupervised Domain Adaptation (UDA) in sparse TAL, which we refer to as Semantic Adversarial unsupervised Domain Adaptation (SADA). Our contributions are threefold: (1) we pioneer the development of a domain adaptation model that operates on realistic sparse action detection benchmarks; (2) we tackle the limitations of global-distribution alignment techniques by introducing a novel adversarial loss that is sensitive to local class distributions, ensuring finer-grained adaptation; and (3) we present a novel set of benchmarks based on EpicKitchens100 and CharadesEgo, that evaluate multiple domain shifts in a comprehensive manner. Our experiments indicate that SADA improves the adaptation across domains when compared to fully supervised state-of-the-art and alternative UDA methods, attaining a performance boost of up to mAP. The code is publicly available at https://github.com/davidpujol/SADA.
1 Introduction
Recent advances in the field of video understanding have played a critical role in the surge of novel video-based applications – e.g., video indexing, summarization or recommendation. A critical task of this field is Temporal Action Localization (TAL), which involves identifying actions in a video consisting of both their time intervals and action categories. This is particularly difficult given the inherent variabilities of videos. These can be presented, among others, in the form of appearance variability – e.g., different kitchens and/or lighting conditions –, acquisition variability – e.g., different recording devices – or viewpoint variability – e.g., first- or third-person. All these prompt a certain degree of confusion between similar actions to be discriminated.
@@ 1 Introduction
Unsupervised Domain Adaptation (UDA) has recently become a hot topic given its potential to leverage unlabelled data to mitigate this domain-induced degradation [18, 40, 23]. Despite its considerable success in image-based tasks [11] the application of UDA to video understanding remains underexplored. In fact, to the best of our knowledge, no prior work addresses UDA for TAL setups. The closest proposal is SSTDA [8], which approaches the problem of action segmentation. This focuses on making per-frame action predictions (i.e., dense TAL), tackling datasets where no concurrent actions take place [16, 29]. This approach inevitably requires additional smoothing techniques to preserve temporal coherence. This motivates the focus of this paper on the more general sparse TAL problem –i.e., segment-level predictions – avoiding the need for additional losses, while intrinsically adapting to multi-label scenarios.
@@
Consequently, in this paper, we propose the first UDA method for sparse multi-label detection on TAL, which we name Semantic Adversarial unsupervised Domain Adaptation, or SADA for short. Concretely, our proposal specifically builds upon an anchor-based architecture given their recent success on sparse TAL [63, 48]. Thus, our goal is to minimize the discrepancy between anchor representations of a labeled source domain and an unlabelled target domain. These anchors are extracted with a multi-resolution architecture [48] that we couple with a novel adversarial loss that improves the limitations of existing UDA works. Concretely, existing works normally align domain distributions globally [18], applying adversarial methods on the feature embeddings regardless of the action class they represent. As we will show, the coarse alignment of anchors of different action classes with even background (no-action) anchors can hurt performance. We propose instead to first use pseudo-labeling [30] to assign an action or background class to each anchor representation. With this, we can factorize the global alignment loss [18] into independent per-class and background distribution alignments. This results in a more sensitive alignment strategy, less prone to semantic feature misalignment – i.e., semantically meaningless alignments across domains – and as we will show, better performing.
@@
Assessing the effectiveness of UDA methods for video understanding is a challenging, still unresolved task. Existing proposals on action segmentation [8] follow a subject-based strategy where they aim to adapt a model to new unseen subjects. Here we refer to subject as a person appearing in a video. Nevertheless, little data is normally available from a single subject, which inevitably requires grouping several of them for training. This allows the model to generalize over the subject variability under study, making it unsuitable for domain adaptation. To address these limitations, we draw inspiration from the work of [64] on action recognition and investigate the impact of viewpoint domain shifts on sparse TAL in CharadesEgo [50]. However, we contend that a more comprehensive evaluation necessitates setups with more controllable shifts. For this, we also propose a suite of 6 new setups based on EpicKitchens100 [14] which study the effect appearance and acquisition domain shifts. These benchmarks demonstrate that SADA mitigates the performance degradation, improving by a large margin the existing fully supervised (namely source-only) and UDA-based proposals. In short, our main contributions are:
@@
We propose for the first time an UDA method suitable for sparse detection scenarios on TAL.
@@ 2 Related work
Temporal Action Localization. At the time of this writing, most of the literature on the task of Temporal Action Localization follows a traditional source-only approach. In other words, they restrict the models’ visibility solely to a training domain, while other domains seen during testing are not available. These works can be categorized as follows: (1) Anchor-based methods [3, 6, 31, 32, 45, 52, 63, 48] propose a two-stage pipeline, consisting of a proposal generation and classification. The first applies heuristic methods – e.g., uniform sampling [3, 63] or action boundaries’ grouping [67, 66] – to generate a dense set of proposals –i.e., temporal segments. In the second stage, they leverage a learnable classifier to predict the corresponding action class and localization offsets of every anchor. Our work falls into this category motivated by the recent success of these methods achieving state-of-the-art results in many TAL benchmarks [25, 65, 14]. (2) Anchor-free methods [49, 33, 32, 62] avoid this two-stage approach making per-frame predictions of their corresponding action labels. These methods, however, often suffer from a tendency towards over-segmentation given the potential discrepancy between neighboring frames. Consequently, they require often complex smoothing techniques to improve the boundary predictions [8]. (3) Query-based methods [53, 36, 35] recently emerged as an alternative paradigm that follows the principles presented by [4]. This approach exploits the use of a Transformer encoder-decoder architecture [56] to learn a fixed small set of queries given refined video features, each identifying one potential action segment. Intuitively, this results in a non-heuristic-based proposal generation. This comes with the limitation of an increased rigidity, as the number of proposals needs to be fixed beforehand.
@@
Unsupervised Domain Adaptation. Domain Adaptation techniques emerge as an effective solution to bridge the gap between data collected from a source and a target distribution, respectively. A large suite of approaches has been proposed to perform this alignment between labeled and unlabeled domains – e.g., discrepancy minimization [28, 60] or entropy minimization [21, 61]. Arguably, nowadays the most popular approach is based on adversarial training [18, 55, 20, 24, 23]. These learn domain-invariant embeddings [12] by training in a min-max fashion a domain classifier to discern if samples come from the source or the target domain. Despite convenient, the simplicity of these methods often degrade the quality of the alignment [30], as they potentially align embeddings of source and target domain that represent different semantic information – e.g., different class labels. Few works have been proposed to do this alignment in a more sensitive way [30]. Works like [59, 24], for instance, reduce the distance of per-class centroids, normally computed as the mean feature embeddings of a given class. Its effectiveness, however, relies on the assumption that the data is distributed somewhat homogeneously around the center, as otherwise, the centroids are not necessarily meaningful. In our work, we couple the advantages of both adversarial domain adaptation and semantic alignment and propose for the first time a pure adversarial semantic loss that yields domain invariant representations in a semantically meaningful way, without making explicit assumptions of the distributions (see Fig. 1).
@@
Unsupervised Domain Adaptation for TAL. Despite the considerable success of UDA methods, their applicability has been mostly restricted to image-based scenarios such as image classification [22, 18, 37, 40] or object detection [9, 42]. Much less attention has been dedicated to video-based applications such as action recognition [7, 26, 43] or spatio-temporal action segmentation [1, 39]. To the best of our knowledge, at the time of this writing, there is no direct comparison with our work focusing on UDA for sparse TAL. The closest work is SSTDA [8] that applies UDA for Action Segmentation. SSTDA proposes the use of two global-distribution-based auxiliary tasks to jointly align cross-domain feature spaces. Unlike our proposal, their work falls into the category of anchor-free, making per-frame action predictions. This restricts its applicability to action segmentation scenarios, where current datasets [16, 29] are designed to deal with frame-based single-action classification. In our work, we overcome this limitation by leveraging an anchor-based architecture that enables a natural adaptation to more realistic multi-label scenarios.
@@ 3.1 Problem definition and notation
In this paper, we address the problem of unsupervised domain adaptation for TAL. For this, we define a source domain and a target domain . Domain consists of labeled input videos , where each video is a sequence of frames with . Here contains the begin, end, and class actions of all the ground-truth (GT) segments of the k-th video, respectively. The target domain is similar to but lacks the GT information. Concretely, it consists of unlabeled input videos . Our goal is to train a model that can identify the action segments, including both segment coordinates and action labels, in videos from domain , while minimizing the performance degradation on the unlabeled domain .
@@ 3.2 Framework overview
We propose a model based on a feature pyramid and an anchor-based classification and localization head (see Fig. 2). This architecture is coupled with a novel semantic adversarial loss that aligns the anchor embeddings across domains and in a semantically meaningful way. More in detail, the model takes as input two videos from domain and , respectively. The model first processes the raw input videos using a frozen pre-trained video backbone. The resulting embeddings of both domains are then passed through a shared SGP pyramid [48] that outputs a set of multi-resolution anchor embeddings for each of the predefined resolution levels. The main goal of our model is to make these anchor embeddings domain invariant. For this, we introduce a level-wise semantic adversarial loss that learns in an adversarial manner to align the embeddings of both domains belonging to a given action class at a given resolution level . Recall that GT information is only available for domain , therefore we rely on the use of pseudo labeling techniques [30] to infer the probable class labels of the data from domain . Finally, we use the domain invariant anchors of domain to train a classification and localization head that learns the underlying tasks in a standard supervised fashion. In short, this permits to learn a classification and localization head that minimizes the decline of performance when applied to the unseen domain .
@@ 3.3 Backbone and SGP pyramid
Our model first takes two input videos and of both domains and . For simplicity, both videos and have length , which we enforce using padding. The method then processes the two videos applying a frozen pre-trained backbone – e.g., I3D [5] or Slowfast [17]. This permits to extract, in an effective way, temporal cues of the video into a set of refined video features. These embeddings are then fed to an SGP feature pyramid [48] which combines the use of SGP blocks and the progressive downsampling of the temporal length by a ratio of . This outputs a set of multi-resolution anchor embeddings and , for the two domains, respectively. Here denotes the set of predefined resolution levels and the anchor embeddings of level of a given domain. Concretely, this permits to obtain embeddings for a set of uniformly sampled anchors at each of the resolution levels. The use of a multi-resolution model is favorable for naturally adapting to different action lengths and abstraction levels.
@@ 3.4 Classification and localization head
To learn the underlying TAL task, we train in a fully supervised manner a classification and localization module with the labeled source domain . Due to the anchor-based nature of our model, we first require a matching strategy between the set of candidate anchors to the actual GT segments. For this, we follow a center sampling strategy [54, 48]. In other words, for a given level , we define an anchor as action anchor if the time instant that it represents is near the center of an action. All the rest are marked as background anchors. We define and as the begins, ends and action classes of their matching GT segments. We identify background anchors with action label 0. With this, we design a classification head that maps each of the anchor embeddings to their class distribution. More specifically, we model this as a sequence of 1D convolutions, and train it using a sigmoid focal loss[34]:
Similarly, we model a localization head identically as , which predicts the begin-end offsets. We thus define the localization loss as a standard mean squared error (MSE) loss over the action anchors only:
where refers to the filtered action-anchor representations from the GT of the -th level only, and to the concatenation operation. This yields the final task loss defined as:
@@ 3.5 Our proposal: Semantic adversarial multi-resolution alignment
One of the main contributions of this paper is the design of a novel adversarial-based loss that we name SADA loss, which attempts to overcome the limitations of the extensively used global adversarial loss [18]. Traditional adversarial domain adaptation relies on the idea of designing a domain classifier that learns to identify the domain that each of the embeddings belongs to. The rest of the model learns concurrently the opposite objective which results in the learning of domain invariant representations[18]. While this approach has been shown to be effective in other fields – e.g., image classification or object detection – we find that its performance in more challenging video understanding setups like TAL presents important challenges. One of the main issues, as argued by [30], is that this loss often suffers from feature misalignment which greatly declines its effectiveness. This refers to the cases where these methods align embeddings of non-matching class labels – i.e., aligning embeddings of domain of an action with embeddings of domain of a class . This issue is further exacerbated in TAL given the noise induced by the alignment of the many background anchors with the action anchors. We study this phenomenon in more depth in the Supp.
@@
Local adversarial alignment: To fix this feature misalignment in realistic scenarios like TAL, we propose an alternative adversarial loss formulation that provides a finer-grained alignment. This loss first attempts to perform a local class-aware alignment. This is, for every given resolution level , we group the action anchors – those matched with a GT action – of and according to their action label . This is straightforward for domain as we have the GT information. In contrast, for domain , we use a hard-pseudo labeling strategy [30] that classifies a given embedding as class if this is the highest-confidence score of the predicted class distribution, and this is above a threshold . Formally we define the pseudo-label of an anchor as:
@@
where are the predicted class probabilities of the anchors. Notice that we mark with class 0 the background anchors, which are not assigned to any action class. From this, we obtain the newly grouped embeddings of source and target domain of class i on level l
@@
for , . Also, is the GT action label of anchor and is its computed pseudo-label from Eq. 4. We then adversarially train a single domain classifier to identify the domain of each of these embeddings using a binary cross entropy (BCE) loss:
@@
where and are the domain labels. We then introduce a Reverse Gradient Layer (GRL)[18] before the discriminator to invert the gradients sign, creating a min-max game where the feature extractor learns to confuse the discriminator. We condition the discriminator to class using a learnable class embedding that we replicate for every selected anchor into an embedding (see Supp. for an ablation of this model decision).
@@
Local and global alignment (SADA): Eq. 7 aims solely to align the action anchors – which are classified as one of the classes – but what happens with the background embeddings that fall below the threshold ? In this case, the loss ignores their influence, yielding only partial alignment.
@@
To overcome this issue, we propose our final SADA loss which attempts to combine the best of both global alignment loss [18] and Eq. 7. For this, we introduce a new loss term for the background anchors as follows:
@@
where again and are the selected background anchors, and is the learnable background embedding. Coupling Eq. 7 and Eq. 8 yields the final formulation of our proposed loss, combining local (class-wise) alignment with the background anchors alignment. Formally:
@@
where is a hyper-parameter that modulates the importance of level on the final alignment loss. See Supp. for an analysis of our model’s sensitivity to this parameter choice.